OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2 (Points : 1) y 9 = 2(x 5) y 5 = 2(x + 9) y + 9 = 2(x 5) y 9 = 2(x + 5)

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

can you help me correct my asnwers

OpenStudy (anonymous):

this onme i am comfident

OpenStudy (anonymous):

but i need help on this one

OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, 7); Slope: (Points : 1) y + 7 =(x 1) y 7 =(x + 1) y + 7 =(x + 1) y 7 =(x 1)

OpenStudy (anonymous):

Is it A

OpenStudy (anonymous):

y+7=-2/3(x-1)

OpenStudy (skullpatrol):

You are missing symbols in your equations???

OpenStudy (anonymous):

its -2/3

OpenStudy (anonymous):

and a and d has minus sighns with the 1 and the x

OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, 7); Slope: (Points : 1) y + 7 =(x 1) y 7 =(x + 1) y + 7 =(x + 1) y 7 =-2/3(x -1)

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

Al of them has -2/3 next to the equal sign

OpenStudy (skullpatrol):

Still missing symbols.

OpenStudy (anonymous):

and with x-1

OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, 7); Slope: (Points : 1) y + 7 =-2/3(x - 1) y - 7 =-2/3(x + 1) y + 7 =-2/3(x + 1) y - 7 =-2/3(x - 1)

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

Fixed it

OpenStudy (anonymous):

is it A the first one

OpenStudy (skullpatrol):

Much better, thank you.

OpenStudy (skullpatrol):

Sub in the point (1, 7) into each of the equations to see which one gives you a true statement: y + 7 =-2/3(x - 1) y - 7 =-2/3(x + 1) y + 7 =-2/3(x + 1) y - 7 =-2/3(x - 1)

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

Number five

OpenStudy (anonymous):

I thnik its

OpenStudy (anonymous):

y=0.012x+.65

OpenStudy (skullpatrol):

y + 7 =-2/3(x - 1) -------> 7+7 =?= (-2/3)(1-1) --> 14 =/=0 y - 7 =-2/3(x + 1)----->7-7 =?= (-2/3)(1+1) ---> 0=/= -4/3 y + 7 =-2/3(x + 1)----->14 =/ -4/3 y - 7 =-2/3(x - 1)-----> 0=0 TRUE!!! <-----ANSWER

OpenStudy (skullpatrol):

I think you are right with y=0.012x+.65